OpenStudy (anonymous):

|2x-4|-1=5 please show work.

OpenStudy (anonymous):

oh i remember this ^^ i did it last week! ok so isolate the absolute value first so get rid of the -1 now what do you have?

OpenStudy (anonymous):

Alright, you're going to want to isolate the absolute value first. So how would you do that?

OpenStudy (anonymous):

2x-4=5?

OpenStudy (anonymous):

yes you can also have 2x-4=-5 or did you learn that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

tay, now just combine like terms and divide, what do you get?

OpenStudy (anonymous):

Not quite, you're going to want to keep those absolute bars up. So first, you're going to want to isolate it. |2x-4|-1=5 Let's get the 1 away from the absolute value bars. You can do that by adding 1 to both sides. What do you get?

OpenStudy (anonymous):

oh yeah 1+5 is 6 so instead of the or fact i said it will be 2x-4=6 or 2x-4=-6

OpenStudy (anonymous):

wait there is another person saying something different

OpenStudy (anonymous):

im confused :(

OpenStudy (anonymous):

sorry i wasn't looking at the problem she is right, add one to both sides then take away the bars

OpenStudy (anonymous):

|2x-4|-1 + 1=5 + 1 |2x-4| = ?

OpenStudy (anonymous):

The 1s on the left side cancel out, so you have: |2x-4| = 5+1 Or |2x-4|=6 Okay, so in order to get rid of the absolute value parts, you have to consider both the positive and negative case: The first case is: 2x - 4 = 6 What's the second case?

OpenStudy (anonymous):

... i don't know

OpenStudy (anonymous):

honey.... just make the = s negative number..... does that make sense?

OpenStudy (anonymous):

|2x-4|=6 could equal either 2x-4=6 (which is the same equation with the bars removed) OR it could be: 2x-4 = -6 (this is the negative case) So you have two equations you have to solve: 2x-4=6 AND 2x-4 = -6

OpenStudy (anonymous):

I'm just helping my brother with his homework

OpenStudy (anonymous):

and i can't explain it so he can understand

OpenStudy (anonymous):

oh just explain it to him by telling him that you have to remember to subtract and divide..... it's really simple so idk what else to say

OpenStudy (anonymous):

i am the brother and i still don't understand you guys

OpenStudy (anonymous):

ok where are you getting confused hun?

OpenStudy (anonymous):

so the problem is 2x-4-1=5 and everyone keeps saying you have to get rid of the 1?

OpenStudy (anonymous):

Yes, because you want to isolate the absolute value (get it alone.)

OpenStudy (anonymous):

ok then what

OpenStudy (anonymous):

how

OpenStudy (anonymous):

so you add the 1 to the 5

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

So now you have: |2x-4|=6 In order to get rid of the absolute value bars, you have to consider both the positive and negative case: The positive case is: 2x - 4 = 6 What's the negative case?

OpenStudy (anonymous):

ok next?

OpenStudy (anonymous):

could you just show me step by step

OpenStudy (anonymous):

Okay, so you have: |2x-4|=6 It's an absolute value equation and in order to solve it, you have to drop the absolute value bars. In order to do that, you have to write two equations (and solve both.) The first case is the easiest to figure out since all you do is drop the bars: 2x - 4 = 6 The second case is where you have to solve the negative case. Here you just make the 6 negative (if it were an inequality you'd have to switch the sign too.) So the second case is: 2x - 4 = -6 So now you have two problems you have to solve: 2x - 4 = 6 AND 2x - 4 = -6

OpenStudy (anonymous):

thank u i think i get it now